OpenStudy (anonymous):

at fixed temperature, the volume of a sample of gas varies inversely as the pressure of the gas. if the pressure of a sample of gas at a fixed temperature is increased by 50% , by what percent is the volume decreased?

OpenStudy (anonymous):

here is given options too...

OpenStudy (anonymous):

and in option there is no such value 3/2 here is just 50%

OpenStudy (anonymous):

i'l check again

OpenStudy (anonymous):

may i give u options...i guess it will be easy...

OpenStudy (anonymous):

a. 25% b. 33*1/3% c. 50% d. 75% e. 100%

OpenStudy (anonymous):

I got 33%

OpenStudy (anonymous):

suppose intial pressure of gas is 100 then v = k/100 it is now increased to 150 so v2 = k / 150 so v2 / v = k /150 / k / 100 = k/150 * 100/k = 2/3 or 67 %

OpenStudy (anonymous):

sry in sol. i did it inc actually its decrease

OpenStudy (anonymous):

thank u soo much both of you...thanks alot...:D

OpenStudy (anonymous):

r u sure jimmyrep? will u check my soln

OpenStudy (anonymous):

the volume is 67% of its initial volume so the decrease is 33%

OpenStudy (anonymous):

oh ok thanks.. you r superhero :)

OpenStudy (anonymous):

are u sure its just 33% cause here is no such option just 33%

OpenStudy (anonymous):

lol ty

OpenStudy (anonymous):

it might be 33*1/3%

OpenStudy (anonymous):

yeah i'm sure its 33 atleast now

OpenStudy (anonymous):

thank u soo much...but as it is question of SAT it is quite tricky.....but m sure the anser is option b. tht is 33%

OpenStudy (anonymous):

thank u so much...its quite tricky cause its question of SAT...but m sure answer is 33%